# Generalizable Whole Slide Image Classification with Fine-Grained Visual-Semantic Interaction

### 29 Feb 2024 v1

@@ 3.3.3 Instance Aggregator Module
We employ self-attention to enable feature interaction among instance features $I_{i}=[e_{i1},e_{i2},\cdots,e_{ij}]$, resulting in the feature $s_{i}$. Subsequently, utilizing the diagnosis prompts $Q$ to aggregate the instance features and acquire the feature $z_{i}$. Then we concatenate $s_{i}$ and $z_{i}$, utilizing the learnable parameter $W$ to fuse these features, ultimately yielding the bag-level feature $v_{i}$. The formulas are shown as follows:
@@ 3.5 Encoder and Loss Function
We divide each WSI into instances $x_{k}$ and encode these instances into embeddings $e_{k}\in R^{D}$ using pre-trained vision encoder $E_{img}$, composed with ResNet structure following [15]. Then we send the instance embeddings into the TFS Module to aggregate the instance features and prompts, and obtain the bag-level embeddings $v_{i}\in R^{D}$. The formulas are shown as follows:
@@
$v_{i}$ $=IA(I_{i},E_{txt}(Q))$ (7)
Besides, we generate pathologically meaningful text embeddings, represented as $t_{i}\in R^{D}$, by leveraging the fine-tuned text encoder BioClinicalBERT [30].
$t_{i}^{c}=E_{txt}(x_{txt})$ (8)
@@ 4.1 Datasets
TCGA Lung Cancer. The TCGA Lung Cancer dataset comprises two cancer subtypes: Lung Adenocarcinoma (LUAD) and Lung Squamous Cell Carcinoma (LUSC). It includes diagnostic slides with 541 slides from 478 LUAD cases and 512 slides from 478 LUSC cases. For WSI preprocessing, follow [15], we cropped each WSI into 256 × 256 non-overlapping patches and removed the background region. The dataset encompasses approximately 5.2 million patches at 20× magnification, averaging about 5,000 patches per WSI. Following [28], we randomly split the dataset into training, validation, and testing sets with a ratio of 65:10:25 on the patient level and adopted 4-fold cross-validation. We collected corresponding pathological diagnosis reports  [^2] https://github.com/tatonetti-lab/tcga-path-reports and cleaned them by GPT-4 to produce fine-grained text description labels. To ensure professionalism and accuracy, we invited professional pathologists to check and correct the textual labels.
Additionally, to evaluate the generalizability of our model and perform zero-shot classification, we curated a dataset of histological subtype labels for TCGA-LUAD from the cBioPortal database [^3] https://www.cbioportal.org/, including 8 histological subtypes of Lung Adenocarcinoma Mixed Subtype, Lung Bronchioloalveolar Carcinoma Nonmucinous, Lung Papillary Adenocarcinoma, and so on. For more detailed information, please refer to the Supplementary Material.
Camelyon16. The Camelyon16 dataset [2] consists of 399 Hematoxylin and Eosin (H&E) stained slide images, utilized for metastasis detection in breast cancer. We preprocessed each WSI by segmenting it into 256 × 256 non-overlapping patches, excluding background regions. In total, this process yields approximately 2.8 million patches at a 20× magnification level, averaging about 7,200 patches in a Bag. We also adopt 3-times 3-fold cross-validation on this dataset.
@@ 4.2 Implementation Details
In experiments, we employed ResNet-18 following [15] as image encoder to extract image features, while pre-trained BioClinicalBERT from [30] as text encoder to generate text features. LoRA [7] was adopted for fine-tuning the text encoder, with an alpha value of 32 and a rank value of 8. We divided the whole slide image into patches with $256\times 256$, then applied a random crop with size $224\times 224$. We adopt AdamW [19] with beta (0.9, 0.98) and eps 1e-8 as our optimizer. Other hyperparameters are learning rate 3e-6, batch size 1 with accumulation step 8, weight decay 1e-4, number of epochs 150, and learning rate warmup ratio 0.1. We utilized mixed-precision training on 4 NVIDIA-A800 GPU.
@@ 4.3 Zero-Shot Histological Subtype Classification
Since only the LUAD’s histological subtype dataset was provided on the online database, we conducted zero-shot subtype classification evaluation on LUAD subtype datasets, with the model being pre-trained on TCGA-LUAD or TCGA-LUSC. Throughout the training phase, subtype label information is deliberately excluded, ensuring that all experiments are conducted solely with fine-grained labels. This aims to evaluate the model’s ability to identify novel diagnostic categories without specific training on subtypes. We extended the morphological appearance text description labels of the target data using GPT-4. After the text encoder encodes the labels, similarity calculation is performed with image features to accomplish zero-shot classification.
@@
Table 2: Zero-Shot performance on histological subtype classification. We pre-train the model using fine-grained labels of TCGA-LUAD and TCGA-LUSC, then apply zero-shot classification to the histological subtypes of TCGA-LUAD.
@@ 4.4 Few-Shot Classification
Our model demonstrates adaptability to various tasks even in scenarios with limited data availability. Few-shot experiments were conducted to demonstrate its transferability to downstream tasks. We initialized the networks with pre-trained weights derived from a model trained on TCGA image-report pairs, and subsequently fine-tuned the model on downstream datasets for few-shot image classification. We followed  [23] and conducted experiments with 1, 2, 4, 8, 16, and additional 0 shots on the downstream dataset. The results are summarized in  Tab. 3.
@@
Our models exhibit remarkable performance in zero-shot classification, achieving an accuracy of 71.26%, surpassing the SOTA method’s one-shot experiment. Upon the introduction of training data, our models display exceptional transferability, outperforming the SOTA by 12.90% in the one-shot setting. The model’s performance improves accordingly with the number of shots. Upon reaching 16 shots, our model reaches an impressive accuracy of 91.25%, showcasing a notable 9.19% improvement close to the fully supervised performance level.
@@ 4.5 Performance Comparison with Existing Works
We compare FiVE with ABMIL [10], DSMIL [15], CLAM-SB [20], CLAM-MB [20], TransMIL [26], DTFD-MIL [36], and MHIM-MIL [28], all of which are attention-based MIL methods. In addition, we included two traditional MIL pooling operations, Max-pooling and Mean-pooling, for comparison. The results of all other methods are reproduced using the official code they provide under the same settings.
To evaluate the performance of FiVE, we conduct fine-grained pre-training exclusively on the TCGA dataset, followed by fully supervised experiments on Camelyon16 and TCGA Lung Cancer datasets for WSI classification. Note that the test data is not used in pre-training. Results are shown in  Tab. 4. It can be found that our method outperforms all the other baselines by a great margin, which fully demonstrates the significance of our fine-grained training scheme in improving performance on downstream tasks. Besides, since the pre-training data mainly comes from TCGA data, on the other hand, the task-specific prompt design is more suitable for TCGA data. This results in the performance improvement of our method for TCGA data being greater than Camelyon16 data.
@@
Table 5: The ablation experiment of the pre-trained model fine-tuned on the TCGA Lung Cancer. SA, FGL, FGG, and LDP represent Self Attention, Fine-Grained Labels, Fine-Grained Guidance, and Learnable Diagnosis Prompts, respectively.
@@ 4.6.1 Effectiveness of TFS Module
We focus on evaluating the impact of the TFS module on the model’s overall performance with the TCGA-Lung dataset. The results detailed in  Tab. 6 reveal significant improvements at each stage of feature enhancement. Initially, the model with only Self Attention attained ACC of 89.77%, AUC of 92.85%, and F1-score of 89.95%. The incorporation of fine-grained labels led to increases of 1.44% in ACC, 1.51% in AUC, and 1.43% in F1-score. Subsequent integration of additional fine-grained guidance further improved performance. Ultimately, our full framework with the TFS Module achieved the highest performance of 94.62% ACC, 96.33% AUC, and 93.89% F1-score. The stepwise addition of these features clearly enhanced the model’s efficacy, underscoring the benefits of combining these techniques in the WSI classification task.
@@ 4.6.2 Effectiveness of Patch Sample Strategy
We verified the impact of different sampling strategies on model performance on TCGA data. Here, we assume that there are two main indicators that affect model performance, sample ratio and max sample threshold (represented as MAXN). At the same time, in order to differentiate the experimental results, we use the unfrozen image encoder for experimental verification.
Figure 3: Classification performance on TCGA Lung Cancer with different sampling strategies, including the average values and the standard deviation (std) values of ACC.
@@ 5 Conclusion
In this paper, we introduce FiVE, a novel framework that demonstrates robust generalization and strong transferability for WSI classification. Our work pioneers the use of non-standardized pathological reports and corresponding WSIs from public databases to develop VLM. To harness the intricacies and diversity present in these reports, we introduce the Task-specific Fine-grained Semantics (TFS) module. This module reconstructs fine-grained labels and corresponding diagnosis prompts during the training phase, while introducing diagnosis prompts, thereby enhancing the semantic relevance of its features. Furthermore, considering that pathological visual patterns are redundantly distributed across tissue slices, we sample a subset of visual patches during training. Our results demonstrate the robust generalizability and computational efficiency of our proposed framework, which also exhibits strong zero-shot performance and is readily adaptable to other tasks with fine-tuning lightly. Moreover, we observe that as the maximum number of sampled patches increases, the model’s performance consistently improves until it plateaus. We aspire to provide empirical insights and contribute to AI pathology research through our methodology.
@@
This supplementary material includes demonstrations of raw unstandardized pathological reports as presented in Sec. A. Additional insights into the labels generated by GPT-4 are detailed in Sec. B. Further elaboration on the prompts utilized for Text Standardization is provided in Sec. C. Moreover, a comprehensive explanation of the Fine-Grained Guidance Construction is outlined in Sec. D.
Appendix A Pathological Report Example
We present a range of pathological report demonstrations in  Fig. 4, highlighting the extracted information utilizing GPT-4 to generate fine-grained text description label, as depicted in  Tab. 6. The sampled diagrams depict considerable variability in the formats of pathological reports stemming from diverse data sources. After the data extraction by GPT-4, the textual descriptions exhibit a more structured format. Although some inconsistencies in expression persist, our objective differs from multi-label classification, emphasizing the semantic expression within the text. This kind of comprehensive description at the slide level enhances our model’s understanding of semantics, consequently refining the model’s ability to generalize.
@@ Appendix A Pathological Report Example
Figure 4: Pathological report examples. We randomly sampled several pathological reports with different reporting standards for display. Sensitive information has been masked.
Table 6: Fine-grained text description labels extracted from raw pathological reports.
@@ Appendix B TCGA Label Description
We utilized the prompt “Describe the morphological characteristics of the LABEL in a single sentence in English.” to obtain label descriptions through GPT-4. When utilized, the placeholder tag LABEL is substituted with each specific label in the process.
@@ B.2 TCGA-LUAD Subtype Label
To evaluate the generalizability of our model and perform zero-shot classification, we curated a dataset of histological subtype labels for TCGA-LUAD , including histological subtypes of 54 samples of Lung Adenocarcinoma Mixed Subtype, 10 samples of Lung Bronchioloalveolar Carcinoma Nonmucinous, 5 samples of Lung Acinar Adenocarcinoma, 1 sample of Mucinous Adenocarcinoma, 4 samples of Micropapillary (colloid) Adenocarcinoma, 3 samples of Lung Bronchioloalveolar Carcinoma Mucinous, 2 samples of Lung Micropapillary Adenocarcinoma, 8 samples of Lung Papillary Adenocarcinoma.
@@ Appendix C Prompts Used for Text Standardization via GPT-4
We present the prompt template employed for GPT-4 to extract information from the raw pathological report, wherein the placeholder tag REPORT is substituted with each distinct raw pathological report. Additionally, alongside furnishing essential prompt information, we’ve manually annotated some assistant examples to aid GPT-4 in further standardizing the output data format.
@@ C.2 Assistant Example
Figure 5: Fine-grained guidance construction pipeline
@@ Appendix D Fine-grained Guidance Construction Pipeline
As depicted in Fig. 5, we partitioned the original fine-grained text descriptions into several parts according to manual-designed prompts. Subsequently, we conducted random sampling and removed “Unknown” tags from the answers. Following this, we shuffled the retained pairs and reconstructed answers into a unified and complete sequence. This approach enables our model to undergo a wider range of description transformations, enriching our text corpus data and consequently enhancing the semantic richness of the model’s features.

### 05 Apr 2024 v2

@@ 3.3.3 Instance Aggregator Module
We employ self-attention to enable feature interaction among instance features $I_{i}=[e_{i1},e_{i2},\cdots,e_{ij}]$, resulting in the feature $s_{i}$. Subsequently, utilizing the diagnosis prompts $Q$ to aggregate the instance features and acquire the feature $z_{i}$. Then we concatenate $s_{i}$ and $z_{i}$, utilizing the learnable parameter $W$ to fuse these features, yielding the bag-level feature $v_{i}$. The formulas are shown as follows:
@@ 3.5 Encoder and Loss Function
We divide each WSI into instances $x_{k}$ and encode these instances into embeddings $e_{k}\in R^{D}$ using pre-trained vision encoder $E_{img}$, composed with ResNet structure following [15], here $D$ represents the dimension of the embeddings. Then we send the instance embeddings into the TFS Module to aggregate the instance features and prompts, and obtain the bag-level embeddings $v_{i}\in R^{D}$. The formulas are shown as follows:
@@
$v_{i}$ $=IA(I_{i},Q)$ (7)
Besides, we generate pathologically meaningful text embeddings for each WSI, represented as $t_{i}\in R^{D}$, by leveraging the fine-tuned text encoder BioClinicalBERT [30].
$t_{i}^{c}=E_{txt}(x_{txt}^{c})$ (8)
@@ 4.1 Datasets
TCGA Lung Cancer. The TCGA Lung Cancer dataset comprises two cancer subtypes: Lung Adenocarcinoma (LUAD) and Lung Squamous Cell Carcinoma (LUSC). It includes diagnostic slides with 541 slides from 478 LUAD cases and 512 slides from 478 LUSC cases. For WSI preprocessing, following [15], we cropped each WSI into 256 × 256 non-overlapping patches and removed the background region. The dataset encompasses approximately 5.2 million patches at 20× magnification, averaging about 5,000 patches per WSI. Following [28], we randomly split the dataset into training, validation, and testing sets with a ratio of 65:10:25 on the patient level and adopted 4-fold cross-validation. We collected corresponding pathology reports  [^2] https://github.com/tatonetti-lab/tcga-path-reports and cleaned them by GPT-4 to produce fine-grained text description labels. To ensure professionalism and accuracy, we invited professional pathologists to check and correct the textual labels. Additionally, to evaluate the generalizability of the model and perform zero-shot classification, we utilized a dataset of histological subtype labels for TCGA-LUAD from the cBioPortal database [^3] https://www.cbioportal.org/, More details about subtype labels are provided in Supplementary Material.
Camelyon16. The Camelyon16 dataset [2] consists of 399 Hematoxylin and Eosin (H&E) stained slide images, utilized for metastasis detection in breast cancer. We preprocessed each WSI by segmenting it into 256 × 256 non-overlapping patches, excluding background regions. In total, this process yielded approximately 2.8 million patches at a 20× magnification level, with about 7,200 patches per bag. We adopted 3-times 3-fold cross-validation.
@@ 4.2 Implementation Details
In experiments, we employed ResNet following [15] as image encoder to extract image features, while pre-trained BioClinicalBERT from [30] as text encoder to generate text features. LoRA [7] was adopted for fine-tuning the text encoder, with an alpha value of 32 and a rank value of 8. We divided the whole slide image into patches with $256\times 256$, then applied a random crop with size $224\times 224$. We adopted AdamW [19] with beta (0.9, 0.98), eps 1e-8, learning rate 3e-6, warmup ratio 0.1, weight decay 1e-4 as our optimizer. Additionally, we used the batch size of 1 with an accumulation step of 8 and trained for 150 epochs. We utilized mixed-precision training on 4 NVIDIA-A800 GPUs.
@@ 4.3 Zero-Shot Histological Subtype Classification
Since only the LUAD’s histological subtype dataset was provided on the online database, we conducted zero-shot subtype classification evaluation on LUAD subtype datasets, with the model being pre-trained on TCGA-LUAD or TCGA-LUSC. Throughout the training phase, subtype label information was deliberately excluded, ensuring that all experiments are conducted solely with fine-grained labels. This aims to evaluate the model’s ability to identify novel diagnostic categories without specific training on subtypes. We extended the morphological appearance text description labels of the target data using GPT-4. After the text encoder encodes the labels, similarity calculation with image features achieved zero-shot classification.
@@
Table 2: Zero-Shot performance on histological subtype classification. We pre-trained the model using fine-grained labels of TCGA-LUAD and TCGA-LUSC, then applied zero-shot classification to the histological subtypes of TCGA-LUAD.
@@ 4.4 Few-Shot Classification
Our model demonstrates adaptability to various tasks even in scenarios with limited data availability. Few-shot experiments were conducted to demonstrate its transferability to downstream tasks. We initialized the networks with pre-trained weights derived from the model trained on TCGA image-report pairs, and subsequently fine-tuned the model on downstream datasets for few-shot image classification. We followed  [23] and conducted experiments with 1, 2, 4, 8, 16, and additional 0 shot on the downstream dataset. The results are summarized in  Tab. 3.
@@
Our model exhibites remarkable performance in zero-shot classification, achieving an accuracy of 71.26%, even surpassing the SOTA method’s one-shot experiment. Upon the introduction of training data, our models display exceptional transferability, outperforming the SOTA by 12.90% in the one-shot setting. The model’s performance improves accordingly with the number of shot. Upon reaching 16-shot, our model reaches an impressive accuracy of 91.25%, showcasing a notable 9.19% improvement close to the fully supervised performance level.
@@ 4.5 Performance Comparison with Existing Works
We compared FiVE with ABMIL [10], DSMIL [15], CLAM-SB [20], CLAM-MB [20], TransMIL [26], DTFD-MIL [36], and MHIM-MIL [28], all of which are attention-based MIL methods. In addition, we included two traditional MIL pooling operations, Max-pooling and Mean-pooling, for comparison. The results of all other methods are reproduced using the official code they provide under the same settings.
To evaluate the performance of FiVE, we conducted fine-grained pre-training exclusively on the TCGA dataset, followed by fully supervised experiments on Camelyon16 and TCGA Lung Cancer datasets for WSI classification. Note that the test data is not used in pre-training. Results are shown in  Tab. 4. It can be found that our method outperforms all the other baselines by a great margin, which fully demonstrates the significance of our fine-grained training scheme in improving performance on downstream tasks. Besides, since the pre-training data mainly comes from TCGA data, on the other hand, the task-specific prompt design is more suitable for TCGA data. This results in the performance improvement of our method for TCGA data being greater than Camelyon16 data.
@@
Table 5: The ablation experiments of the pre-trained model fine-tuned on the TCGA Lung Cancer. SA, FGL, FGG, and LDP represent Self Attention, Fine-Grained Labels, Fine-Grained Guidance, and Learnable Diagnosis Prompts, respectively.
@@ 4.6.1 Effectiveness of TFS Module
We focused on evaluating the impact of the TFS module on the model’s overall performance with the TCGA Lung Cancer. The results detailed in  Tab. 5 reveal significant improvements at each stage of feature enhancement. Initially, the model with only Self Attention attained ACC of 89.77%, AUC of 92.85%, and F1-score of 89.95%. The incorporation of fine-grained labels led to increases of 1.44% in ACC, 1.51% in AUC, and 1.43% in F1-score. Subsequent integration of additional fine-grained guidance further improved performance. Ultimately, full framework with the TFS Module achieved the highest performance of 94.62% ACC, 96.33% AUC, and 93.89% F1-score. These ablation experiments highlight the benefits of integrating these methods in the WSI classification task.
@@ 4.6.2 Effectiveness of Patch Sample Strategy
We verified the impact of different sampling strategies on model performance on TCGA data. Here, we assumed that there are two main indicators that affect model performance, sample ratio and max sample threshold (represented as MAXN). At the same time, in order to differentiate the experimental results, we used the unfrozen image encoder for experimental verification.
Figure 3: Classification performance on TCGA Lung Cancer with diverse sampling strategies, presenting average and standard deviation (std) ACC values.
@@ 5 Conclusion
In this paper, we introduce FiVE, a novel framework that demonstrates robust generalization and strong transferability for WSI classification. Our work pioneers the utilization of non-standardized WSI-report pairs from public databases to develop a VLM. To capture the complexities and diversity within these reports, we introduce the Task-specific Fine-grained Semantics (TFS) module. This module reconstructs fine-grained labels and diagnosis prompts during training, enhancing the semantic relevance of its features by introducing diagnosis prompts. Furthermore, considering that pathological visual patterns are redundantly distributed across tissue slices, we employ a sampling strategy to reduce computational costs. Our experiments demonstrate the robust generalizability and computational efficiency of the proposed framework, which can also be easily adapted to other tasks with minimal fine-tuning. We aspire to provide empirical insights and contribute to AI pathology research.
@@
This supplementary material includes demonstrations of raw unstandardized pathology reports as presented in Sec. A. And the detailed process of generating label descriptions using GPT-4, as well as specific label descriptions, are provided in Section B. Additionally, further elaboration on the prompts utilized for Text Standardization is provided in Sec. C. Moreover, a comprehensive explanation of the Fine-Grained Guidance Construction is outlined in Sec. D.
Appendix A Pathology Report Example
We present a range of pathology report demonstrations in  Fig. 4, highlighting the extracted information utilizing GPT-4 to generate fine-grained text description label, as depicted in  Tab. 6. The sampled diagrams depict considerable variability in the formats of pathology reports stemming from diverse data sources. After the data extraction by GPT-4, the textual descriptions exhibit a more structured format. Although some inconsistencies in expression persist, our objective differs from multi-label classification, emphasizing the semantic expression within the text. This kind of comprehensive description at the slide level enhances our model’s understanding of semantics, consequently refining the model’s ability to generalize.
@@ Appendix A Pathology Report Example
Figure 4: Pathology report examples. We randomly sample several pathology reports with different reporting standards for display. Sensitive information has been masked.
@@
Table 6: Fine-grained text description labels extracted from raw pathology reports.
@@ Appendix B TCGA Label Description
We utilize the prompt “Describe the morphological characteristics of the LABEL in a single sentence in English." to obtain label descriptions through GPT-4. When utilized, the placeholder tag LABEL is substituted with each specific label in the process.
@@ B.2 TCGA-LUAD Subtype Label
To evaluate the generalizability of our model and perform zero-shot classification, we curate a dataset of histological subtype labels for TCGA-LUAD , including histological subtypes of 54 samples of Lung Adenocarcinoma Mixed Subtype, 10 samples of Lung Bronchioloalveolar Carcinoma Nonmucinous, 5 samples of Lung Acinar Adenocarcinoma, 1 sample of Mucinous Adenocarcinoma, 4 samples of Micropapillary (colloid) Adenocarcinoma, 3 samples of Lung Bronchioloalveolar Carcinoma Mucinous, 2 samples of Lung Micropapillary Adenocarcinoma, 8 samples of Lung Papillary Adenocarcinoma.
@@ Appendix C Prompts Used for Text Standardization via GPT-4
We present the prompt template employed for GPT-4 to extract information from the raw pathology report, wherein the placeholder tag REPORT is substituted with each distinct raw pathology report. Additionally, besides providing essential prompt information, we have manually annotated some assistant examples to assist GPT-4 in further standardizing the output data format.
@@ C.2 Assistant Example
Figure 5: Fine-grained guidance construction pipeline.
@@ Appendix D Fine-grained Guidance Construction Pipeline
As depicted in Fig. 5, we partition the original fine-grained text descriptions into several parts based on manually designed prompts. Next, we perform random sampling and remove “Unknown” tags from the answers. Afterward, we shuffle the retained pairs and reconstruct the answers into a unified and complete sequence.. This approach enables our model to undergo a wider range of description transformations, with changes in granularity providing diverse perspectives on the visual image. Aligning visual images with text descriptions of varying granularities enriches the training data, offering more descriptive perspectives and semantic information for visual patterns. This enhances the semantic richness of the model’s features and facilitates model transfer to downstream tasks.